OpenStudy (anonymous):

A test consists of 20 questions, and students are asked to only answer 15 of them. In how many different ways can they choose 15 questions?

OpenStudy (anonymous):

So to find how many ways to choose some number of things from another number you do the permutation of them only if the order doesn't mattert: \[P(20,15)\] \[P(n,k)=\frac{ n! }{ k!(n-k!) }\] So do: \[\frac{ 20! }{ 15!(20-15!) }\] Simplify:\[\frac{ 20! }{15!* 5! }\] So the answer is 20!/15!*5! Whatever that is just type it in on google or a calculator :

OpenStudy (anonymous):

it should give you: 15504 ways to arrange the questions and do them in any order you want

OpenStudy (anonymous):

btw \[P(n,k) \] in my answer up there should be:\[\left(\begin{matrix}n \\ k\end{matrix}\right)\]

OpenStudy (tkhunny):

\(20\choose 15\)?